# NonGEMM Bench: Understanding the Performance Horizon of the Latest ML Workloads with NonGEMM Workloads

### v3

@@ Abstract
Among ML operators today, GEneralMatrix Multiplication (GEMM)-based operators are known to be key operators that construct the main backbone of ML models. As their computational overhead dominates the overall execution time (e.g., 49% - 94%in our results), GEMM-based operators have been the prime optimization target for fast ML inference and training, which led to advanced GPUs and accelerators available today. Such GEMM accelerators boost the GEMM performance and provide significant end-to-end speedup compared to CPUs, aligned with the lesson from Amdahl’s law. However, accelerating GEMM has shifted the amdahl’s law’s landscape for ML inference; due to the decreased GEMM execution time, the relative execution time of non-GEMM operators is not dominant. Nonetheless, the performance of non-GEMM operators has not been studied with the latest GPUs. In addition, recent models include more complex non-GEMM operators (e.g. SiLU), which motivates the characterization of the performance to guide the future ML system design.
Therefore, we first extensively profile the execution time and energy of 18 widely adopted ML models in Hugging Face and Torchvision on mobile, workstation, and data center platforms with/without GPUs. Based on the results, we analyze the type of non-GEMM operators that contribute most to the execution time for each task. Finally, we collect all models with our performance profiling harness and create a new ML benchmark, NonGEMM Bench, which automatically reports detailed breakdown into GEMM and NonGEMM operators. NonGEMM Bench also provides a collection of 1460 non-GEMM operators with realistic input arguments and data to guide the non-GEMM operator-oriented system optimizations. Our case study shows the non-GEMM operators account for 60% of the total execution time, on average, across theree hardware platforms with GPUs. In addition, we report that the dominant non-GEMM operator varies with the task and even for different ML deployment software tools. Our results indicate that considering non-GEMM operator is important for future hardware and software system designs for ML systems.
1 Introduction
@@
Figure 1: The latency breakdown into GEMM and non-GEMM operators on AMD EPYC 7763 + NVIDIA A100 GPU. We measure the latency on two popular models from HuggingFace (a) GPT2-XL (batch 1) [5] and (b) Vision Transformer (batch 1) [25].
However, because the GEMM operators are being accelerated, the non-GEMM operators, such as memory operations (e.g., reshape, view, and transpose), normalization, and logit computation functions (e.g., Softmax), now account for a considerable amount of the end-to-end latency, compared to that of GEMM operators.  Figure 1 shows the profiled latency breakdown into GEMM and non-GEMM operators running inferences on state-of-the-art large language (GPT2-XL [5]) and image classification (Vision Transformer [25]) models. The motivational data imply a major shift of the landscape of Amdahl’s law [21] for ML acceleration, indicating that we now need to consider non-GEMM operators in the ML system optimization. However, no study thoroughly explored the non-GEMM operator performance in latest models, although it is important to quantitatively understand the current status due to the fast evolving GEMM accelerators such as GPUs.
Therefore, we propose NonGEMM Bench, which focuses on non-GEMM operator analysis in popular ML models today, and perform a thorough peformance characterization of Non-GEMM operators in the latest models widely adopted by the community. Unlike previous benchmarks, NonGEMM Bench thoroughly analyzes NonGEMM operator performance, providing important insights on the ML system optimization with powerful accelerators. NonGEMM Bench consists of 18 models collected on Hugging Face [43] and Torchvision [30] on various computer vision and natural language processing tasks.
To enable non-GEMM-oriented system optimization, NonGEMM Bench provides a microbenchmark of non-GEMM operators. In addition to the preset of non-GEMM operators collected in popular Hugging Face models, NonGEMM Bench can profile arbitrary non-GEMM operators supported by PyTorch FX  [37] and ONNX  [8], which provides flexibility to users. The flow allows users to analyze any non-GEMM operators without manually writing custom functions, unlike previous benchmark [27]. In addition, NonGEMM Bench utilizes input argument specification extracted from real data. This helps analyze non-GEMM operators in realistic settings.
Using this benchmark, we perform case studies on 18 popular Hugging Face and Torchvision models on image classification, object detection, segmentation, and language models (e.g., Llama 2 [41] and GPT-2 [35]). We study the performance on different hardware configuration ranging from mobile, workstation, to data center, which includes four CPU and four GPU models. In our case studies, we evaluate the effect of GPU acceleration on the relative latency across operators of NonGEMM Bench models and highlight that the contribution of non-GEMM operators to the end-to-end latency of the accelerated model.
We will open-source NonGEMM Bench at the time of publication to facilitate the ML system research aligned with the current performance horizon. Including that, we make the following contributions:
@@
  • We propose NonGEMM Bench, which provides analysis on non-GEMM operators in the 18 latest popular models from Hugging Face and Torchvision.
  • To facilitate non-GEMM operator-aware optimizations, we organize a microbenchmark of non-GEMM operators collected in the popular models and provide realistic input arguments collected from realistic input data.
  • Using NonGEMM Bench, we perform case studies on three different hardware configurations ranging from mobile, workstation, and to data center, and show the non-GEMM operators are becoming major consideration across all platforms.
  • We identify different set of dominant non-GEMM operators depending on the task domain and deployment SW flow, which motivates domain-oriented specialization and shows the importance of deployment software’s role for non-GEMM operators.
2 Background
@@
To introduce NonGEMM Bench, we discuss the ML operators and recent ML models.
2.1 ML Operators
ML operators are the building blocks for ML models, which define the computation over one or multiple input tensors. Examples include convolution (Conv2d), matrix multiplication (linear, BMM, etc.), activation, and normalization functions, as listed in Figure 2. Operators can be categorized into two classes: GEMM operation-based and the others, which we term as non-GEMM operators. We discuss each class of ML operators next.
2.1.1 GEMM-based Operators (GEMM Operators)
@@
GEMM-based operators are known to be compute-intensive, which accounts for the majority of the execution unless accelerated by GPUs or accelerators, as CPU results in in Figure 1 show. However, they have regular computation patterns that can be summarized as a perfectly nested for loop. The regular pattern allows various loop optimization techniques such as loop reordering, tiling, and parallelization, which is referred to as dataflow [33, 26, 46]
With the dominance of GEMM operators in execution time and high optimization potential together, GEMM operators have been the prime optimization target for acceleration, which led to high-performance GPUs (e.g., H100 [14]) and accelerators [23].
2.1.2 Non-GEMM Operators
@@
To understand the extent of the non-GEMM operators, we analyze non-GEMM operators in 18 recent models in the computer vision and natural language processing domains. We select models based on their popularity in the Hugging Face community to obtain realistic workload. We list the models we investigated in Table 1. Based on our analysis, we categorize non-GEMM operators based on their functionality and summarize their usage in models and characteristics in Table 2.
  • Normalization. Normalization operators regularize the data range across a selected dimension using the mean and standard deviation statistics. Examples include BatchNorm [22] and LayerNorm [7], which are widely adopted in computer vision and NLP ML models [18, 5]. Such variants defer by the choice of the normalized dimension (e.g., BatchNorm normailizes a tensor across the batch dimension). RMS Norm [47], which is adopted in recent large language models [41], is another example of the normalization function. RMS Norm eliminates the division by standard deviation in typical normalization functions and performs $\sqrt{\frac{1}{n}\Sigma^{n}_{i=1}(X_{i}-\mu)}$, where $X_{i}$, n, and $\mu$ refer to the i-th data, number of data, and the mean, respectively.
@@
  • Memory Operators. Memory operators are responsible for the allocation and the layout modification of tensors inside the memory. Examples cover the reshape and concatenation (cat) operators. Reshape operators do not create new tensors. They modify the shape (e.g. dimension order) of a tensor and return a new view of the tensor following the new dimension order. While reshape operators do not allocate additional memory, the cat operator allocates new memory to store a newly created tensors by concatenating the input tensors along a new dimension.
  • Arithmetic. Arithmetic operators cover non-GEMM arithmetic operators applied on tensors in an element-wise or reduction pattern. Element-wise division applied to scale the elements of tensors in the attention block as shown in Figure 3 (c) is an example of such operators.
@@
2.2 ML Models and Popular Tasks
The heterogeneity in non-GEMM operators enabled ML engineers and researchers to build models supporting a wide range of tasks (e.g. Computer Vision and NLP). As highlighted in  Figure 3, computer vision ( Figure 3 a and b) and NLP ( Figure 3) models are characterized by distinct architectures leveraging unique combinations of GEMM and non-GEMM operators.
@@ 2.2 ML Models and Popular Tasks
3 NonGEMM Bench
Understanding the performance landscape of ML models with non-GEMM workloads necessitates a specialized benchmark. A good benchmark targeting non-GEMM operators should 1) capture the operator level performance breakdown not only in a standalone scenario, but also in an end to end inference to understand the effects of the non-GEMM operators. Moreover, it should analyze 2) frequently used models in 3) diverse task domains on 4) real data to reflect the actual performance of realistic application.
@@ 3 NonGEMM Bench
ML Benchmarks available today (e.g., MLPerf [36]), unfortunately, do not satisfy the requirements since they do not focus on the non-GEMM operators. Long-tail bench [27] identified a similar problem as this work, but it focuses on a limited set of custom kernels, which fails to represent broad task domain. To better-understand the impact of non-GEMM operators on system performance, we develop a new ML benchmark, NonGEMM Bench, that satisfies all the requirements for a non-GEMM operator specific benchmark. NonGEMM Bench provides operator-level breakdown of execution time at the computational/operator graph level in end-to-end inference scenarios, which enables detailed non-GEMM operator performance analysis (e.g., impact of each operator type on the end-to-end performance), as we present in Section 4. In addition, to facilitate non-GEMM operator-focused system optimizations, NonGEMM Bench includes a micro-benchmark of non-GEMM operators with concrete input arguments and tensors collected from realistic datasets and models. We discuss the datasets and models adopted in NonGEMM Bench and describe the structure of NonGEMM Bench in detail next.
3.1 NonGEMM Bench Models
@@
To capture the performance in the latest ML workload, we select 18 highly downloaded models from HuggingFace [1] to enhance the representativeness of NonGEMM Bench.  Table 1 lists the NonGEMM Bench model registry (Figure 4) which contains 18 models based on state-of-the-art CNN and Transformer architectures with number of parameters ranging from 0.85M to 7B, demonstrating the diverse model coverage of NonGEMM Bench. The selected models cover four major task domains in ML, which include Image Classification (IC), Object Detection (OD), Image Segmentation (IS), and Natural Language Processing (NLP).
3.1.1 Computer Vision (CV)
@@
Deep Learning is widely used in the CVdomain [25]  [19]  [38]. NonGEMM Bench covers three important tasks of the domain:
@@
Image Classification (IC). Image classification refers to a CV task that identifies a class label of a given image. Image classification models extract features (i.e., high-level and -dimensional information of the input image) from an input image and reports the class label utilizing the features. NonGEMM Bench includes eight most popular IC models in HuggingFace [1], which includes ResNet50  [19], MobileNetv2  [39], three variants of Vision Transformer [25], and three variants of Swin Transformer  [29].
@@
Image Segmentation (IS). Image segmentation refers to a computer vision task that identifies the area in a image for each class. Like IC models, IS models also extract features and utilize them to identify objects located in an image and spatially separate them by highlighting pixels belong to each object. NonGEMM Bench includes two state-of-the-art IS models: Segformer  [45] and MaskFormer  [12]. Object Detection (OD). Object detection refers to a computer vision task that identifies the location of objects in an image and outputs the bounding box of each object. Like aforementioned models, OD models extract features. In addition to that, OD models generate region proposal, which refers to the candidate location and bounding boxes for objects in an image. Using Region of Interest (RoI) processing non-GEMM operators listed in Table 2, OD models refine the raw region proposal generated by a region proposal network. The refined RoIs are used as inputs to the classifier CNN at the end, and the classifier determines the class label of object inside each refined RoI. NonGEMM Bench includes three popular OD models [2]: FasterRCNN  [38], MaskRCNN  [18], and DETR  [10].
3.1.2 Natural Language Processing (NLP)
@@
NLP refers to tasks involving the analysis and understanding of human (natural) language. NLP models extract context and features from an input text sequence and use the extracted context and features to perform multiple applications like translation and text generation [48] [9]. Transformer [42] based DNNs have become the dominant model architecture in NLP and are the backbone of popular state of the art generative large language models like GPT  [35] and Llama  [41].  Figure 3 (c) shows the layer architecture of GPT’s transformer. It consists of a self attention block build by cascading GEMM operators with Normalization, Memory, Logit Comuputation and Element-wise Arithmetic non-GEMM oeprators ( Table 2). NonGEMM Bench includes five popular NLP models  [3]: Bert [24], three variants of GPT2 [5], and Llama2-7b [41].
3.2 Benchmark Suite
NonGEMM Bench provides two flows: end-to-end inference and microbench flows. The End-to-end flow evaluates the performance of operators when the model is running inference on dataset data. The Microbench flow performs micro-benchmarking on the operator by running it in a standalone fashion using synthetic data with realistic dimensions. We discuss input, output, and the benchmark software structure for each flow.
3.2.1 NonGEMM Bench Inputs
@@
NonGEMM Bench receives target deployment software flow and other configurations as inputs. Optionally, users can provide custom models with associated datasets to profile their workloads. By default, NonGEMM Bench runs the 18 models in Table 1. We describe each input in detail as follows:
@@
Models. As described in  Section 3.1, NonGEMM Bench offers a registry of pre-selected popular ML models for users to analyze. Nonetheless, we designed NonGEMM Bench to be easily expandable to accommodate rapidly evolving ML models that constantly introduce new operators. Users can benefit from the features of our benchmark by simply plugging their new models into the NonGEMM Bench model registry (Figure 4) by specifying the model class, its weights and any data preprocessor.
@@
Deployment Flow. NonGEMM Bench deploys its models using two popular inference frameworks: ONNX Runtime [16] and PyTorch [34] eager mode, TorchScript, and TorchDynamo introduced in PyTorch 2.0 [6].
@@
Datasets. To evaluate the models, NonGEMM Bench utilizes real data from popular datasets in each domain. We use ImageNet 2012  [15] and MS COCO  [28] for computer vision tasks. As for language models, we use wikitext dataset [31] available on HuggingFace [4].
@@
Misc. Configurations. This includes all other configurations for profiling runs, which includes the batch size, the number of profiling iterations, and the target device.
Table 1: Tasks and Models in NonGEMM Bench
@@
Image Classification ResNet50 [19] 0.85M ImageNet [15]
MobileNetv2 [39] 3.4M
ViT large 16 [25] 307M
ViT huge 14 [25] 632M
Swin tiny [29] 29M
Swin small [29] 50M
Swin base [29] 88M
Object Detection FasterRCNN [38] 42M COCO [28]
MaskRCNN [18] 44M
@@
Segmentation Maskformer [12] 102M COCO [28]
SegFormer [45] 3.7M
Language Models GPT2 [5] 117M wikitext [31]
GPT2 Large [5] 762M
GPT2 X-Large [5] 1.5B

### v4

# Understanding the Performance Horizon of the Latest ML Workloads with NonGEMM Workloads

@@ Abstract
Among ML operators today, GEneralMatrix Multiplication (GEMM)-based operators are known to be key operators that build the main backbone of ML models. As their computational overhead dominates the overall execution time (e.g., 42.8% - 96.6% in our results), GEMM operators have been the prime optimization targets for fast ML inference. This led to advanced GPUs and accelerators available today, which provided significant boost in the GEMM performance compared to CPUs, aligned with the lesson from Amdahl’s law. However, accelerating GEMM has significantly shifted the Amdahl’s law’s landscape for ML inference; due to the decreased GEMM execution time, the relative execution time of non-GEMM operators is not dominant. Although the importance of non-GEMM performance is increasing, we have little knowledge about the non-GEMM performance horizon in the latest hardware platforms and models.
Therefore, to guide non-GEMM-oriented optimizations, we conduct a thorough performance analysis of 16 widely adopted ML models in Hugging Face and Torchvision on workstation and data center platforms with/without GPUs. We discover that non-GEMM performance bottleneck is a considerable issue across all the platforms and models, accounting for 11.3% to 73.6% of total latency, on average. The challenge significantly aggravates when we apply quantization, which is a common model compression technique, due to the boosted GEMM performance and extra non-GEMM operators for dequantization and requantization. To provide insights into non-GEMM optimization targets, we demystify the most dominant non-GEMM operators for each model and deployment software. We also show that widely adopted optimizations such as operator fusion do not completely address the non-GEMM performance bottleneck, where non-GEMM operators still account for 15% to 48% of total latency. We will open-source our non-GEMM-oriented benchmark framework to facilitate research in non-GEMM optimization.
I Introduction
@@
Figure 1: The latency breakdown into GEMM and non-GEMM operators on AMD EPYC 7763 + NVIDIA A100 GPU. We measure the latency on two popular models from HuggingFace (a) GPT2-XL (batch 1) [4] and (b) Swin Transformer (batch 1) [37].
However, because the GEMM operators are being accelerated, the non-GEMM operators, such as memory operations (e.g., reshape, view, and transpose), normalization, and logit computation functions (e.g., Softmax), now account for a considerable amount of the end-to-end latency, compared to that of GEMM operators.  Figure 1 shows the profiled latency breakdown into GEMM and non-GEMM operators running inferences on state-of-the-art large language (GPT2-XL [4]) and image classification (Swin Transformer [37]) models. The motivational data imply a major shift of the landscape of Amdahl’s law [24] for ML inference acceleration, indicating that we now need to consider non-GEMM operators as one of the major optimization targets in the ML system optimization. However, the research community today lacks a thorough and systematic performance analysis and characterization of non-GEMM operators in the latest models, which hinders the development of non-GEMM oriented optimization techniques.
Therefore, we collect widely-used ML models from Hugging Face [61] and Torchvision [38] and perform a thorough performance characterization of of non-GEMM operators in the 18 latest models of four major task domains: Image Classification (IC), Image Segmentation (IS), Object Detection (OD), and Natural Language Processing (NLP). We evaluate the effect of GPU acceleration on the relative latency across GEMM and non-GEMM operators in collected models and conduct deep-dive analysis on the impact of different hardware platforms (mobile, workstation, and datacenter), deployment software, and common optimizations (operator fusion and quantization). Based on our case studies, we highlight that the non-GEMM performance challenge is common in accelerated inferences and existing optimization techniques (e.g., operator fusion) cannot completely address the challenges. Also, we demystify the most time-consuming non-GEMM operators in each model, which will help the research community identify non-GEMM operators to be optimized.
To facilitate such research in non-GEMM-oriented optimization techniques, we build an open-source benchmark specialized in non-GEMM performance anlysis, NonGEMM Bench, which will be released after publication. NonGEMM Bench can profile arbitrary non-GEMM operators supported by PyTorch [46], ONNX [7], and TensorRT [43], in addition to the preset of non-GEMM operators collected from the selected 16 popular models, which provides desired flexibility to users for follow-up research.
We summarize our contributions as follows:
@@
  • We perform case studies on two different hardware configurations, workstation and data center, and show the non-GEMM operators are becoming a major consideration across all platforms.
  • We identify different dominant non-GEMM operators depending on the model and deployment software flow, which indicates that non-GEMM optimization need to be specialized for each model and deployment software.
  • We analyze the impact of a common non-GEMM-aware optimization, operator fusion and show that operator fusion does not completely mitigate non-GEMM bottleneck for all models, which motivates follow-up research in non-GEMM performance optimizations
  • We evaluate the performance of non-GEMM operators with quantization and show the non-GEMM bottleneck aggravates with data.
  • We will open-source NonGEMM Bench, an expandible benchmark flow that enables thorough non-GEMM performance characterization for any model supported by ONNX runtime, Tensor RT, and PyTorch, to facilitate non-GEMM-oriented research.
II Background
@@
II-A ML Operators
ML operators are the building blocks for ML models, which define the computation over one or multiple input tensors. Examples include convolution (Conv2d), matrix multiplication (linear, BMM, etc.), activation, and normalization, as listed in Figure 2. We categorize operators into two classes: GEMM operation-based and the others, which we term as non-GEMM operators. We discuss each class of ML operators next.
II-A1 GEMM-based Operators (GEMM Operators)
@@
GEMM-based operators are known to be compute-intensive, which accounts for the majority of the execution unless accelerated by GPUs or accelerators, as CPU results in in Figure 1 show. However, they have regular computation patterns that can be summarized as a perfectly nested for loop. The regular pattern allows various loop optimization techniques such as loop reordering, tiling, and parallelization, which is referred to as dataflow [45, 33, 66] With the dominance of GEMM operators in execution time and high optimization potential together, GEMM operators have been the prime optimization target for acceleration, which led to high-performance GPUs (e.g., H100 [13]) and accelerators [29].
II-A2 Non-GEMM Operators
@@
To understand the extent of the non-GEMM operators, we analyze non-GEMM operators in 16 recent models in the computer vision and natural language processing domains. We select models based on their popularity in the Hugging Face to obtain realistic workload. We list the models we investigated in Table I. Based on our analysis, we categorize non-GEMM operators based on their functionality and summarize their usage in models and characteristics in Table II.
  • Normalization. Normalization operators regularize the data range across a selected dimension using the mean and standard deviation. Examples include BatchNorm [26] and LayerNorm [6], which are widely adopted in computer vision and NLP models [21, 4]. RMS Norm [67], which is adopted in recent large language models [58], is another example of the normalization function. RMS Norm eliminates the division by standard deviation in typical normalization functions and performs $\sqrt{\frac{1}{n}\Sigma^{n}_{i=1}(X_{i}-\mu)}$, where $X_{i}$, n, and $\mu$ refer to the i-th data, number of data, and the mean, respectively.
@@
  • Memory Operators. Memory operators are responsible for the memory allocation and the layout modification of tensors. For example, Reshape modifies the shape (e.g. dimension order) of a tensor and return a new view of the tensor following the new dimension order.
  • Element-Wise Arithmetic. Elemennt-wise arithmetic operators cover non-GEMM operators applied on individual elements in a tensor. For example, Figure 3 (c) contains an element-wise division applied to scale the elements of tensors in the attention block
@@
II-B ML Models and Popular Tasks
The heterogeneity in non-GEMM operators enabled ML developers to build models supporting a wide range of tasks (e.g. Computer Vision and NLP). As highlighted in  Figure 3, computer vision ( Figure 3 a and b) and NLP ( Figure 3) models are characterized by distinct architectures leveraging unique combinations of GEMM and non-GEMM operators.
@@ II-B ML Models and Popular Tasks
III Performance Characterization Methodology
To understand the realistic performance landscape of the latest ML models with non-GEMM workloads, we must (1) capture operator level performances in end-to-end inferences to understand, (2) use widely-used models by the research community and industry, (3) cover diverse task domains, and (4) use real datasets.
@@ III Performance Characterization Methodology
However, ML Benchmarks available today (e.g., MLPerf [52]), unfortunately, do not satisfy all the requirements since they do not focus on the non-GEMM operators. Long-tail bench [34] identified a similar problem as this work, but it focuses on a limited set of custom kernels, which fails to represent broad task domain. Therefore, to facilitate our non-GEMM operators analysis and better-understand the impact of non-GEMM operators on system performance, we develop a new ML benchmark, NonGEMM Bench. NonGEMM Bench provides operator-level breakdown of end-to-end inference latency in the operator graph level, which enables detailed non-GEMM operator performance analysis, as we present in Section IV. To capture the performance in the latest ML workload, we select 15 highly downloaded (more than 10K downloads on average) models from HuggingFace [1] to enhance the representativeness of NonGEMM Bench and our analysis. We discuss the datasets and models in NonGEMM Bench in detail and describe the structure of NonGEMM Bench next.
III-A Evaluated Models
@@
Table I lists the NonGEMM Bench model registry which contains 15 models based on state-of-the-art CNN and Transformer architectures with number of parameters ranging from 0.85M to 7B, demonstrating the diverse model coverage of NonGEMM Bench. The selected models cover four major task domains in ML, which include Image Classification (IC), Object Detection (OD), Image Segmentation (IS), and Natural Language Processing (NLP).
III-A1 Computer Vision (CV)
@@
Deep learning is widely used in the CV domain [32]  [22]  [53]. NonGEMM Bench covers three important tasks in the domain:
@@
Image Classification (IC). Image classification refers to a CV task that identifies a class label of a given image. Image classification models extract features (i.e., high-level and dimensional information of the input image) from an input image and report the class label utilizing the features. NonGEMM Bench includes eight most popular IC models in HuggingFace [1]: Three variants of Vision Transformer [32], and three variants of Swin Transformer  [37].
@@
Image Segmentation (IS). Image segmentation refers to a computer vision task that identifies the area in a image for each class. Like IC models, IS models also extract features and utilize them to identify objects located in an image and spatially separate them by highlighting pixels that belong to each object. NonGEMM Bench includes two state-of-the-art IS models: Segformer  [64] and MaskFormer  [10]. Object Detection (OD). Object detection refers to a computer vision task that identifies the location of objects in an image and outputs the bounding box of each object. OD models extract features and generate region proposals, which refer to the candidate locations and bounding boxes of objects. Using Region of Interest (RoI) processing non-GEMM operators in Table II, OD models refine raw region proposals generated by a region proposal network. The refined RoIs are used as inputs to the CNN classifier at the end, and the classifier determines the class label of objects inside each refined RoI. NonGEMM Bench includes three popular OD models [2]: FasterRCNN  [53], MaskRCNN  [21], and DETR  [9].
III-A2 Natural Language Processing (NLP)
@@
NLP refers to tasks involving the analysis and understanding of human (natural) language. NLP models extract context and features from an input text sequence and use the extracted context and features to perform multiple applications like translation and text generation [68] [8]. Transformer [59] based DNNs have become the dominant model architecture in NLP and are the backbone of popular state-of-the-art generative large language models like GPT  [50] and Llama  [58].  Figure 3 (c) shows the layer architecture of GPT’s transformer. It consists of a self-attention block built by cascading GEMM operators with Normalization, Memory, Logit Comuputation and Element-wise Arithmetic non-GEMM oeprators ( Table II). NonGEMM Bench includes five popular NLP models  [3]: Bert [31], three variants of GPT2 [4], and Llama2-7b [58].
III-B Profiling Methodology
To enable the thorough analysis of the landscape of non-GEMM performance presented in Section IV, NonGEMM Bench provides an end-to-end profiling flow to breakdown the inference latency at the operator level. We discuss input, output, and the benchmark software structure.
III-B1 NonGEMM Bench Inputs
@@
NonGEMM Bench receives target deployment software flow (PyTorch, ONNX runtime, or TensorRT), dataset, and other configurations listed in Figure 4 as inputs. Optionally, users can provide custom models with associated datasets to profile their workloads. By default, NonGEMM Bench runs the 15 models in Table I. We describe each input in detail as follows.
@@
Models. As described in  Section III-A, NonGEMM Bench includes a registry of 15 selected popular ML models. Nonetheless, we designed NonGEMM Bench to be easily expandable to accommodate rapidly evolving ML models that constantly introduce new operators. Users can benefit from the features of our benchmark by simply plugging their new models into the NonGEMM Bench model registry (Figure 4) by specifying the model class, its weights and any data preprocessor.
@@
Deployment Flow. NonGEMM Bench deploys its models using three popular inference frameworks: ONNX Runtime [17], PyTorch [46] TensorRT [43].
@@
Datasets. To evaluate the models, NonGEMM Bench utilizes real data from popular datasets in each domain. We use ImageNet 2012  [15] and MS COCO  [36] for computer vision tasks. As for language models, we use wikitext dataset [39] available on HuggingFace. For custom models, users can specify their own dataset for the model.
@@
Misc. Configurations. This includes the batch size, the number of profiling iterations, and the target device.
TABLE I: Tasks and Models Evaluated in Section IV
@@
Image Classification (IC) ViT base (Vt-b) [32] 307M ImageNet  [15]
ViT large (Vt-l) [32] 307M
ViT huge (Vt-h) [32] 632M
Swin tiny (Sw-t) [37] 29M
Swin small (Sw-s) [37] 50M
Swin base (Sw-b) [37] 88M
Object Detection (OD) FasterRCNN (FRCNN) [53] 42M COCO  [36]
MaskRCNN (MRCNN) [21] 44M
@@
Image Segmentation (IS) Maskformer (MF) [10] 102M COCO  [36]
SegFormer (Seg) [64] 3.7M
Natural Language Processing (NLP) GPT2 [4] 117M wikitext  [39]
GPT2 Large (gpt2-l) [4] 762M
GPT2 X-Large (gpt2-xl) [4] 1.5B